# Study of silicon photomultipliers for the readout of a lead/scintillating-fiber calorimeter

### v1

@@ 5.1 Consistency check on the measurement method
Some checks were made during the process, in order to validate the performed measurements. The first consistency check is shown in Fig. 15, where the amplitude of signals taken from two separate channels of the same SiPM matrix are compared. As expected, the distribution of events is well described by a bisector. This result confirms the uniformity of the photon collection event-by-event on the SiPM array. This uniformity was already observed on statistical basis (see Fig. 5, right).
@@ 6 Conclusions
A piece of the KLOE calorimeter has been equipped with SiPMs and PMTs in order to compare their performance for a potential application in an element (SAND) of the Near Detector complex of the future DUNE experiment. The PMT coupling was optimized in the KLOE calorimeter and it was kept unchanged for this study. Various optical couplings have been tested for adapting the SiPMs with the KLOE light guides and it was verified that the partial coverage of the light guide does not prevent reaching a proper readout of the signal. Efficiency and timing resolution have been measured and compared for both detectors. It was found that the SiPM noise cannot be neglected in the efficiency estimate. Specifically, the noise rejection implies a non-negligible reduction of the possible SiPM efficiency making it lower by a few percent with respect to the PMT ones. For the timing resolution, it was observed that PMTs reach a value lower than 220 ps for the scintillation signal, while SiPMs exhibit a resolution above 240 ps. Even if the difference is small, PMTs seem to perform better.
In conclusion, the difficulties in coupling SiPMs to ECAL without deep mechanical changes, the lack of improvement, the cost, and the necessary commissioning time, advise against the substitution of the 4880 available and tested PMTs with new SiPMs in the KLOE calorimeter to be reused in the Near Detector complex in DUNE. Nonetheless, the results from this study do not exclude the possible use of SiPMs for other calorimetry applications.
@@ Acknowledgments
We would like to thank Stefano Miscetti (INFN, Laboratori Nazionali di Frascati) who delivered to us the stack of lead foils and scintillating fibers. We thank also Antonio Di Domenico (Sapienza, Rome) and Alessandro Montanari (INFN, Sezione di Bologna) for their fruitful comments and suggestions.

### v2

@@ 5.1 Consistency check on the measurement method
Some checks were made during the process, in order to validate the performed measurements. The first consistency check is shown in Fig. 15, where the amplitude of signals taken from two separate channels of the same SiPM matrix are compared. As expected, the distribution of events is well described by a bisector. This result confirms the uniformity of the photon collection event-by-event on the SiPM array. This uniformity was already observed on a statistical basis (see Fig. 5, right).
@@ 6 Conclusions
A cut-out of the KLOE calorimeter has been equipped with SiPMs and PMTs in order to compare their performance for a potential application in an element (SAND) of the Near Detector complex of the future DUNE experiment. The PMT coupling was optimized in the KLOE calorimeter and it was kept unchanged for this study. Various optical couplings have been tested for adapting the SiPMs with the KLOE light guides and it was verified that the partial coverage of the light guide does not prevent reaching a proper readout of the signal. Efficiency and timing resolution have been measured and compared for both detectors. It was found that the SiPM noise cannot be neglected in the efficiency estimate. Specifically, the noise rejection implies a non-negligible reduction of the possible SiPM efficiency making it lower by a few percent with respect to that of the PMTs. For the timing resolution, it was observed that PMTs reach a value lower than 220 ps for the scintillation signal, while SiPMs exhibit a resolution above 240 ps. Even if the difference is small, PMTs seem to perform better.
In conclusion, the difficulties in coupling SiPMs to ECAL without major mechanical changes, the lack of improvement, the cost, and the necessary commissioning time, advise against the substitution of the 4880 available and tested PMTs with new SiPMs in the KLOE calorimeter to be reused in the Near Detector complex in DUNE. Nonetheless, the results from this study do not exclude the possible use of SiPMs for other calorimetry applications.
@@ Acknowledgments
We would like to thank Stefano Miscetti (INFN, Laboratori Nazionali di Frascati) who delivered us the stack of lead foils and scintillating fibers. We thank also Antonio Di Domenico (Sapienza, Rome) and Alessandro Montanari (INFN, Sezione di Bologna) for their fruitful comments and suggestions.